OpenStudy (anonymous):

If 5x+y=19 and x-3y=7, then x+y=?

OpenStudy (anonymous):

5x+y=19, set x equal to zero. then solve. x-3y=7, set y equal to zero. then solve. then plug into x+y=?

OpenStudy (anonymous):

Do you know how to do as a system of equations problem

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Could you show me how to do it that way

OpenStudy (anonymous):

okay so your equation will look like this: 5x+y=19 x-3y=7 then you will multiply -5 to the bottom equation. so that the x's will cancel out and all you have to solve is for y. once you solve for y plug that answer to either problem and you will get your answer

OpenStudy (anonymous):

Okay thank you i get it now

OpenStudy (anonymous):

your welcome